Circular No. 8697

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN CYGNUS S. Nakano, Sumoto, Japan, reports the discovery by H. Nishimura

(Miyawaki, Kakegawa, Japan) of a possible nova (mag 10.5) on two Kodak T-Max 400 films taken on Apr. 2.807 UT using a Pentax 6×7 camera (+ 200-mm f/4.0 lens; 60-s exposures). Nishimura provided the following position for the new object: α = 21^{h}11^{m}32^{s}, δ = +44^{o}48′02′′ (equinox 2000.0), noting that nothing was present at this location on his films taken on Mar. 28 (limiting mag 12) or on earlier patrol films back to Oct. 2001. Following a request to the AAVSO, R. Miles (Stourton Caundle, Dorset, England) reports a confirmation of the object at V = 8.5 ± 0.3 on a CCD image taken on Apr. 4.995, providing the following precise position end figures: 32^{s}.34, 03′′.9 (uncertainty estimated as ± 0′′.3). Miles states that nothing is visible at this location on a Digitized Sky Survey image. H. Yamaoka, Kyushu University, notes that a very red star of mag ∼ 14 in several star catalogues has position end figures ∼ 31^{s}.9, 03′′.3.

SUPERNOVAE 2006bg, 2006bh, 2006bi, 2006bj Four additional apparent supernova have been reported from unfiltered

CCD images: 2006bg and 2006bi by R. Holmes, Charleston, IL (cf. IAUC 8677; 2006bg with V. Hoette, and 2006bi with H. Devore and P. Miller); 2006bh by L. A. G. Monard (cf. IAUC 8693); and 2006bj by R. Quimby (cf. IAUC 8680).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006bg Mar. 27.28 11^{h}35^{m}25.^{s}17 +54^{o}57′37′′.4 18.4 16′′.5 W, 2′′.3 N
2006bh Apr. 2.12 22 40 16.10 −66 29 06.3 14.8 48′′ W, 22′′ S
2006bi Mar. 25.39 13 59 11.52 +27 52 17.5 18.1 0′′.2 W, 2′′.6 N
2006bj Mar. 26.19 12 22 19.16 + 7 37 25.3 17.3 1′′.1 E, 1′′.1 N

Additional approximate magnitudes for 2006bg: 1993 Apr. 25, [21.0 (Palomar Sky Survey, red); 2006 Jan. 26.38 UT, [20.4 (Holmes); Apr. 2.18, 18.3 (Holmes). Additional magnitudes for 2006bh in NGC 7329: Jan. 17.751, [17.5; Apr. 3.086, 14.6 ± 0.1. Nothing is visible at the location of 2006bh on the Digitized Sky Survey (limiting red mag 20.5). Additional magnitudes for 2006bi: 1996 July 8, [21.0 (PSS, red); 2006 Feb. 19 and 24, [20.2 (Holmes); Mar. 27.31, 18.2 (Holmes). Additional approximate magnitudes for 2006bj: 2005 May 9, [19.0; 2006 Apr. 2.17, 17.2. A spectrum of 2006bj taken on Mar. 3.34 by S. C. Odewahn and F. Deglman shows it to be a young type-II supernova (details on CBET 460).

2006 April 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green